if diameter of circle is increased by 40% ,then its area increases by ? solve it free
1
Expert's answer
2013-01-24T09:28:17-0500
The formula for the area of a circle is S0=(Pi*d^2)/4,where d is the diameter of circle and Pi=3.14, S0 - area of circle before
increasing.
After increasing the diameter will be 1.4d and the area will be :
S1=(Pi*(1.4d)^2)/4=(1.96Pi*d^2)/4=1.96S0
Then the area of circle increased by 96%